OpenStudy (kirbykirby):

Understanding this proof for the proposition "For all integers a, gcd(9a+4, 2a+1) = 1. Proof: gcd(9a+4, 2a+1) = gcd(2a+1, a) = gcd(a, 1). Since gcd(a, 1)=1, gcd(9a+4, 2a+1) =1. I don't understand what they are doing in this proof? How did they find all this?

OpenStudy (anonymous):

first line because 4(2a+1)=8a+4 and 9a+4-(8a+4)= a

OpenStudy (anonymous):

second line because a times 2 =2a and 2a+1-2a=1

OpenStudy (anonymous):

although the second equality is more or less obvious since 2a+1 leaves a remainder of 1 when divided by a

OpenStudy (anonymous):

good morning myininaya how was early am calc?

myininaya (myininaya):

i'm kindof mad my class stood me up i could have slept longer one student showed i thought integration by parts lol

myininaya (myininaya):

taught*

OpenStudy (anonymous):

oh i like that line one student shows up, and you will definitely be having integration by parts part today, part next class, part next week...

myininaya (myininaya):

i was like heres 5 bonus points enjoy

myininaya (myininaya):

i think i will give a quiz on integration by parts tomorrow

OpenStudy (anonymous):

sounds good now if i don't get out of here and finish my work, i am going to be in real trouble

OpenStudy (anonymous):

good idea!

myininaya (myininaya):

lol all of them will fail except that one

OpenStudy (anonymous):

right! it must be annoying. i had a class where about half showed up last week, so we did fun stuff

OpenStudy (anonymous):

i taught them how to add

myininaya (myininaya):

lol integration by parts is fun

OpenStudy (anonymous):

actually i think so too

myininaya (myininaya):

i showed the one student the table i will not show the table anymore hes the only one who gets the easy way

OpenStudy (anonymous):

maybe he will share the secret off to the garage, see you later

myininaya (myininaya):

don't leave me i forgot -1 comes before 0 i need your help

OpenStudy (anonymous):

what problem?

OpenStudy (anonymous):

minus one comes before zero?

myininaya (myininaya):

its the problem before this one i feel like an idiot i was like choose a number before -1 like 0

OpenStudy (anonymous):

there is a limit problem for you to do, x - sinx etc i will look at your problem